A1 (driving gear) has 40 teeth and A2 (driven gear) has 82 teeth. If A1 rotates 10 teeth, how many degrees does A2 rotate?
Answer #1

degree of rotation of A2 / degree of rotation of A1 = 1/Gear ratio = 40/82 __________________ eq.1

=> Degree of rotation of A2 = 10*40/82 = 4.878 degrees ________________________ Answer